# QServe: W4A8KV4 Quantization and System Co-design for Efficient LLM Serving

### v2

@@ VI-C Efficiency Evaluation
Ablation study on quantization techniques
we examine the impact on accuracy of various quantization techniques implemented in QoQ. Our analysis begins with round-to-nearest (RTN) W8A8 quantization on Llama-2-7B (per-channel + per-token). We then lower the quantization precision and apply different techniques step-by-step. For each step, we evaluated the WikiText2 perplexity and end-to-end inference performance on L40S with 64 requests of 1024 input tokens and 512 output tokens. The results are detailed in Figure 16. We see that reducing the weight precision to 4 bits significantly impaired the model performance, though it increased end-to-end processing speed by 1.12$\times$ and saved 3.5GB GPU memory. Rotating the block input modules helped suppress the activation outliers, resulting in 0.18 perplexity improvement. In addition, minimizing the block output MSE through weight clipping further decreased the perplexity by 0.16. Consequently, our W4A8 configuration has achieved a perplexity comparable to that of W4A16. However, quantizing KV cache to 4 bits again deteriorated model performance by 0.14, although it substantially enhanced the end-to-end inference throughput by 1.47$\times$ and halved GPU memory usage. To solve this problem, SmoothAttention reduced perplexity by 0.05, without adding system overhead. Progressive group quantization further improved perplexity by an additional 0.02, with only a negligible increase in dequantization overhead. Lastly, activation-aware channel reordering enhanced perplexity by 0.03.
@@
Dequantization overhead: We measure the dequantization overhead of per-group QServe-W4A8 GEMM and other baselines in Figure 18. Our dequantization overhead is comparable with TRT-LLM-W4A16, but since we perform computation on INT8 tensor cores, we enjoy 2$\times$ higher throughput.
@@
Comparisons under the same batches: We demonstrate speedup results under the same batch sizes in Figure 17. For Llama-2-7B, we show that the 1.88$\times$ speedup over TRT-LLM can be broken down to two parts: 1.45$\times$ from same batch speedup and 1.3$\times$ from the enlarged batch size. For larger models like Llama-2-13B, scaling up the batch size and single batch speedup are equally important (1.7$\times$ improvement).
@@
Improvement breakdown for KV4 attention: We detail the enhancements from attention optimizations in Section Section V-C. Starting with the basic KV4 implementation, which exhibits an A100 latency of 0.48ms for a 64$\times$1024 input, the application of bit tricks from [20] reduces the kernel latency to 0.44ms. Further improvements are achieved by simplifying the control flow, which reduces latency by an additional 0.05ms. Subsequently, converting the QK and SV products to FP16 each contributes a 0.03ms latency reduction. Asynchronous prefetching of dequantization parameters at the start of the attention kernel further enhances performance, ultimately reducing the latency to 0.28ms and achieving an end-to-end improvement of 1.7$\times$.
VII Related Work
@@
Quantization of LLMs. Quantization reduces the size of LLMs and speedup inference. There are two primary quantization strategies: (1) Weight-only quantization [12, 23, 10, 19] benefits edge devices where the workload is memory-bound, improving weight-loading speed. However, for cloud services with high user traffic and required batch processing, this method falls short as it does not accelerate computation in compute-bound scenarios. (2) Weight-activation quantization accelerates computation in batch processing by quantizing both weights and activations [8, 36, 38]. OmniQuant [30] and Atom [44] exploring more aggressive quantizations (W4A4, W4A8) and mixed precision to enhance model quality and efficiency, though these can impact model accuracy and reduce serving throughput. QuaRot [2] further refines W4A4 by rotating weights and activations at the cost of increased computational overhead due to additional transformations required during inference.
@@
LLM serving systems. Numerous systems have been proposed for efficient LLM deployment. Orca [40] employs iteration-level scheduling and selective batching in distributed systems. vLLM [22] features virtual memory-inspired PagedAttention, optimizing KV cache management. SGLang [45] enhances LLM programming with advanced primitives and RadixAttention. LMDeploy [7] offers persistent batching and blocked KV cache features to improve deployment efficiency. LightLLM [6] manages GPU memory with token-wise KV cache control via Token Attention, increasing throughput. MLC-LLM [32] utilizes compiler acceleration for versatile LLM deployment across edge devices. TensorRT-LLM [25] is the leading industry solution and the most important baseline in this paper.
@@
LLM Accelerators. Transformers and LLMs have also generated considerable research interest in domain-specific accelerator design. Several works, such as $A^{3}$ [14], ELSA [15], and SpAtten [35], have applied pruning techniques to the attention module, while GOBO [41] and EdgeBERT [31] have investigated quantization approaches. Additionally, DOTA [27] introduces a lightweight, runtime detector for omitting weak attention connections, coupled with specialized accelerators for transformer inference. Apart from attention optimizations, STA [11] leverages $N$:$M$ sparsity and specialized softmax module to reduce off-chip communication. Moreover, DFX [16] exploits model parallelism and optimized dataflow for low-latency generation. However, these accelerators have yet to be scaled up to recent LLMs with billions of parameters.
VIII Conclusion

### v3

@@ 6.2 Accuracy Evaluation
6.3 Efficiency Evaluation
We present relative performance comparisons in Figure 15 and absolute throughput values in Table 4. We use per-channel quantization for A100 and per-group quantization for L40S. This is because L40S has stronger CUDA cores for dequantization. Relative to the best-performing configuration of TensorRT-LLM, QServe demonstrates significant improvements on A100: it achieves 2$\times$ higher throughput for Llama-1-30B, 1.2-1.4$\times$ higher throughput for Llama-2 models, 1.2$\times$ higher throughput for Mistral and Yi, and 2.4$\times$ higher throughput for Qwen-1.5. The performance improvements are particularly notable on the L40S GPUs, where we observed a throughput improvement ranging from 1.47$\times$ to 3.47$\times$ across all seven models evaluated. Remarkably, despite the L40S’s significantly smaller memory capacity compared to the A100, QServe effectively maintains the same batch size as TensorRT-LLM on the A100. This achievement is attributed to our aggressive 4-bit quantization applied to both weights and the KV cache. By examining Table 4, we clearly observe that serving five of seven models under 34B on L40S with QServe achieves even higher throughput than serving them on A100 using TensorRT-LLM. Our performance gain over Atom and QuaRot on A100 is even more prominent since these systems did not outperform TensorRT-LLM. On L40S, QServe still achieves 10% higher throughput than Atom when running Llama-2-7B, the only model supported by their system despite the fact that we use higher quantization precision. Besides, the accuracy achieved by QServe is much better than Atom, as indicated in Table 3.
6.4 Analysis and Discussion.
Ablation study on quantization techniques. We examine the impact on accuracy of various quantization techniques implemented in QoQ. Our analysis begins with round-to-nearest (RTN) W8A8 quantization on Llama-2-7B (per-channel + per-token). We then lower the quantization precision and apply different techniques step-by-step. For each step, we evaluated the WikiText2 perplexity and end-to-end inference performance on L40S with 64 requests of 1024 input tokens and 512 output tokens. The results are detailed in Figure 16. We see that reducing the weight precision to 4 bits significantly impaired the model performance, though it increased end-to-end processing speed by 1.12$\times$ and saved 3.5GB GPU memory. Rotating the block input modules helped suppress the activation outliers, resulting in 0.18 perplexity improvement. In addition, minimizing the block output MSE through weight clipping further decreased the perplexity by 0.16. Consequently, our W4A8 configuration has achieved a perplexity comparable to that of W4A16. However, quantizing KV cache to 4 bits again deteriorated model performance by 0.14, although it substantially enhanced the end-to-end inference throughput by 1.47$\times$ and halved GPU memory usage. SmoothAttention reduced perplexity by 0.05, without adding system overhead. Progressive group quantization further improved perplexity by an additional 0.04, with only a negligible increase in dequantization overhead. Lastly, activation-aware channel reordering enhanced perplexity by 0.03.
@@ 6.3 Efficiency Evaluation
Ablation study on QServe system: Dequantization overhead. We measure the dequantization overhead of per-group QServe-W4A8 GEMM and other baselines in Figure 18. Our dequantization overhead is comparable with TRT-LLM-W4A16, but since we perform computation on INT8 tensor cores, we enjoy 2$\times$ higher throughput.
@@
Comparisons under the same batches. We demonstrate speedup results under the same batch sizes in Figure 17. For Llama-2-7B, we show that the 1.88$\times$ speedup over TRT-LLM can be broken down to two parts: 1.45$\times$ from same batch speedup and 1.3$\times$ from the enlarged batch size. For larger models like Llama-2-13B, scaling up the batch size and single batch speedup are equally important (1.7$\times$ improvement).
@@
Improvement breakdown for KV4 attention. We detail the enhancements from attention optimizations in Section Section 5.3. Starting with the basic KV4 implementation, which exhibits an A100 latency of 0.48ms for a 64$\times$1024 input, the application of bit tricks from Kim et al. (2022) reduces the kernel latency to 0.44ms. Further improvements are achieved by simplifying the control flow, which reduces latency by an additional 0.05ms. Subsequently, converting the QK and SV products to FP16 each contributes a 0.03ms latency reduction. Asynchronous prefetching of dequantization parameters at the start of the attention kernel further enhances performance, ultimately reducing the latency to 0.28ms and achieving an end-to-end improvement of 1.7$\times$.
7 Related Work
@@
Quantization of LLMs. Quantization reduces the size of LLMs and speedup inference. There are two primary quantization strategies: (1) Weight-only quantization Frantar et al. (2022); Lin et al. (2024); Dettmers et al. (2023b); Kim et al. (2024) benefits edge devices where the workload is memory-bound, improving weight-loading speed. However, for cloud services with high user traffic and required batch processing, this method falls short as it does not accelerate computation in compute-bound scenarios. (2) Weight-activation quantization accelerates computation in batch processing by quantizing both weights and activations Dettmers et al. (2022); Wei et al. (2022); Xiao et al. (2023). OmniQuant Shao et al. (2023) and Atom Zhao et al. (2023) exploring more aggressive quantizations (W4A4, W4A8) and mixed precision to enhance model quality and efficiency, though these can impact model accuracy and reduce serving throughput. QuaRot Ashkboos et al. (2024) further refines W4A4 by rotating weights and activations at the cost of increased computational overhead due to additional transformations required during inference.
@@
LLM serving systems. Numerous systems have been proposed for efficient LLM deployment. Orca Yu et al. (2022) employs iteration-level scheduling and selective batching in distributed systems. vLLM Kwon et al. (2023b) features virtual memory-inspired PagedAttention, optimizing KV cache management. SGLang Zheng et al. (2023) enhances LLM programming with advanced primitives and RadixAttention. LMDeploy Contributors (2023b) offers persistent batching and blocked KV cache features to improve deployment efficiency. LightLLM Contributors (2023a) manages GPU memory with token-wise KV cache control via Token Attention, increasing throughput. MLC-LLM team (2023) utilizes compiler acceleration for versatile LLM deployment across edge devices. TensorRT-LLM NVIDIA (2023) is the leading industry solution and the most important baseline in this paper.
@@
LLM Accelerators. Transformers and LLMs have also generated considerable research interest in domain-specific accelerator design. Several works, such as $A^{3}$ Ham et al. (2020), ELSA Ham et al. (2021), and SpAtten Wang et al. (2021), have applied pruning techniques to the attention module, while GOBO Zadeh et al. (2020) and EdgeBERT Tambe et al. (2021) have investigated quantization approaches. Additionally, DOTA Qu et al. (2022) introduces a lightweight, runtime detector for omitting weak attention connections, coupled with specialized accelerators for transformer inference. Apart from attention optimizations, STA Fang et al. (2022) leverages $N$:$M$ sparsity and specialized softmax module to reduce off-chip communication. Moreover, DFX Hong et al. (2022) exploits model parallelism and optimized dataflow for low-latency generation. However, these accelerators have yet to be scaled up to recent LLMs with billions of parameters.
8 Conclusion
@@ Acknowledgements
Appendix A Artifact Appendix
A.1 Abstract
This artifact contains necessary scripts and dependencies to faithfully reproduce the crucial experiments presented in the paper. To successfully run the experiments, a host system with x86_64 CPUs is required, along with at least one A100 or L40S NVIDIA GPU. We also provide a pre-built docker image to simplify the environment setup process.
A.2 Artifact check-list (meta-information)
  • Program: Efficiency benchmarking code for QServe; efficiency benchmarking code for baseline systems such as TensorRT-LLM.
  • Compilation: Completed in the docker.
  • Transformations: N/A.
  • Binary: N/A.
  • Model: Llama-3-8B, Llama-2-7B, Mistral-7B, Llama-2-13B.
  • Data set: None.
  • Run-time environment: NVIDIA Container Toolkit (nvidiadocker).
  • Hardware: A host with x86_64 CPUs and at least one NVIDIA A100 GPU (recommended) or L40S GPU.
  • Run-time state: N/A.
  • Execution: All benchmarks are executed on NVIDIA GPUs, while some data pre-processing code is executed on the host CPU.
  • Metrics: LLM generation throughput.
  • Output: Generation throughput (tokens/second).
  • Experiments: Inference speed measurement for QServe and baseline systems such as TensorRT-LLM.
  • How much disk space required (approximately)?: 512G.
  • How much time is needed to prepare workflow (approximately)?: Around 1 hour to pull docker images depending on the Internet connection and CPU performance.
  • How much time is needed to complete experiments (approximately)?: Around 1 hour to finish the efficiency benchmarks of QServe; and 2-4 GPU hours to finish the TensorRT-LLM benchmarks depending on the GPU performance and number of tasks to evaluate.
  • Publicly available?: Yes.
  • Code licenses (if publicly available)?: Apache License 2.0.
  • Data licenses (if publicly available)?: MIT.
  • Workflow framework used?: Docker.
  • Archived (provide DOI)?: https://doi.org/10.5281/zenodo.14991385
A.3 Description
A.3.1 How delivered
We will provide AE reviewers with a pre-built docker image containing QServe, TensorRT-LLM and all necessary dependencies.
A.3.2 Hardware dependencies
A host machine with x86_64 CPUs and at least one NVIDIA A100 GPU (recommended) or L40S GPU.
A.3.3 Software dependencies
A GPU-compatible Docker runtime environment is required.
A.4 Installation
We recommend that users utilize our pre-built Docker images to set up the environment and run all experiments within the GPU-supported Docker container.
1docker run --gpus all -it --workdir /root shang12138/qserve-mlsys25-ae
A.5 Experiment workflow
The generation throughputs of QServe and baseline system (i.e., TensorRT-LLM) can be measured with the following commands.
1# QServe benchmark
2cd /root/QServe
3bash scripts/benchmark/benchmark_a100.sh
4# Run this command if you run on A100 GPU
5# Results in ./parsed_results.csv
6
7# TensorRT-LLM benchmark
8cd /root/TensorRT-LLM
9bash launch-all.sh
10# Launch TensorRT-LLM evaluation
11# Results in ./results.csv
A.6 Evaluation and expected result
Table 6: Generation throughput of QServe and baseline (TensorRT-LLM). Unit: tokens/second.
Model TensorRT-LLM (W8A8KV8) QServe
Llama-3-8B 2387.55 2980.69
Llama-2-7B 2339.97 2860.01
Mistral-7B 2427.64 3031.93
Table 6 provides reference numbers for QServe benchmarks. Please note that absolute throughput measurements may vary slightly, even on identical GPU platforms, due to differences in machine conditions. However, the relative acceleration ratios should remain consistent.
A.7 Experiment customization
The users are encouraged to carry out experiments with different models and batch sizes by modifying the benchmarking scripts. Accuracy evaluation is omitted to simplify the environment setup. The accuracy results can be reproduced with open-source library deepcompressor.
A.8 Methodology